# Contrastive Visual Data Augmentation

### v1

@@ 6 Conclusion
Acknowledgment
This material is based on research supported by the ECOLE program under Cooperative Agreement HR00112390060, with the US Defense Advanced Research Projects Agency (DARPA). The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing DARPA, or the U.S. Government.
@@ Appendix A Appendix
A.1 Qualitative Comparison
Figure 3: Qualitative Comparison of Visual Data Augmentation Methods. Contrastive textual + visual features are used for CoDA. Phyllobates Samperi is a new species of hypertoxic poison dart frog first described in 2024, currently it does not have a common name, the concept is an example from the NovelSpecies dataset.
Figure 3 illustrates the qualitative comparison of the quality of generated images using CoDA and other approaches.
The Crop method, while useful for localized feature emphasis, often results in the loss of crucial visual details necessary for species identification. For instance, in the Phyllobates Samperi images, cropping removes the black spots on the frog’s skin, which are an essential distinguishing feature. Without these patterns, the cropped images lack key identity cues, potentially leading to misclassification. Similarly, in the Tail-Spot Wrasse images, cropping reduces visibility of the distinct horizontal striping pattern along the fish’s body, making it difficult to recognize key species attributes.
ARMADA is capable of retaining some structural features, but it struggles with precise reproduction due to the limited image editing capabilities. This limitation is particularly evident in its generated images, where critical patterns such as the Phyllobates Samperi’s orange stripes are missing. The generated frog appears to have a distorted pattern, failing to fully capture the contrast between black skin and bright orange lines, which are key species identifiers. Similarly, in the case of the Tail-Spot Wrasse, the image generated by ARMADA loses the feature of its vibrant horizontal stripes, leading to a visually inconsistent and less biologically accurate representation.
In contrast, CoDA successfully captures all species-specific features by leveraging contrastive textual and visual attributes through different image generation models. CoDA (SD-3.5) produces a high-fidelity image of Phyllobates Samperi, accurately preserving the orange stripes, dark skin, and black spots. However, slight variations in texture suggest that this model, while effective, may not fully match the real-world skin reflectivity of the species. Meanwhile, CoDA (Recraft V3) generates an even more realistic image, successfully capturing the frog’s signature features with improved color richness and anatomical precision, making it nearly indistinguishable from real-world references.
Generally, CoDA (SD-3.5) and CoDA (Recraft V3) both perform significantly better than previous methods. Take Tail-Spot Wrasse as another example, the horizontal stripes, which were previously distorted or missing in previous methods, are now clearly visible. CoDA (Recraft V3), in particular, produces a more vivid and structurally accurate representation, ensuring the preservation of both color gradients and fin structure.
Note that the quality of CoDA-generated images is inherently dependent on the capability of the underlying image generation model, meaning that limitations in the base model, such as resolution constraints or texture inaccuracies, may impact the fidelity of the final augmented data.
@@ A.3.3 Image Generation and Verification
Post-generation, we perform automated verification using LLaVA V1.6-34b, running on an A6000 GPU. Each image would takes approximately 1 minutes to run for feature presence using a feature-matching confidence threshold of 0.85. Images with a satisfaction rate $S(i^{\text{synthetic}},\mathcal{F},\mathcal{M})<1.0$ are discarded.

### v2

@@ 7 Conclusion
Acknowledgment
This material is based on research supported by the ECOLE program under Cooperative Agreement HR00112390060 with the US Defense Advanced Research Projects Agency (DARPA), an award from Office of Naval Research with #N00014-23-1-2780, Apple Research Award, and Amazon AGI Research Award. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing DARPA, or the U.S. Government.
@@ Appendix A Appendix
A.1 Limitations and Future Work
Our work is not without limitations. First, in our experiments, we focus on the fine-tuning use case as it is the most general and intuitive way to utilize our augmented visual data. In the future, we plan to investigate other conceivable use cases for our augmented data, include model adaptation (Sung et al., 2022), test-time augmentation (Gidaris & Komodakis, 2018), visual information retrieval (Wu et al., 2025), and more. Third, the modularity of our method also invites other researchers to replace components of CoDA with superior models to achieve better performance. The NovelSpecies dataset, which can be updated with new species in future years, may also be used to evaluate future VLMs’ novel concept recognition abilities. Finally, we also expect improved versions of T2I generation-based visual data augmentation techniques to eventually surpass CoDA in effectiveness and efficiency. Potential improvements may include more robust image / feature filtering and more controllable text-conditioned image generation like multi-view synthesis. We hope our work can pave the way for future downstream advancements by demonstrating effective uses of our augmented visual data for enhancing model capabilities.
@@ A.3.3 Image Generation and Verification
Post-generation, we perform automated verification using LLaVA V1.6-34b, running on an A6000 GPU. Each image would take approximately 1 minutes to run for feature presence using a feature-matching confidence threshold of 0.85. Images with a satisfaction rate $S(i^{\text{synthetic}},\mathcal{F},\mathcal{M})<1.0$ are discarded.